OpenStudy (eunice_o):

Factor the following expression . 16v^8 x^9 -28v^2x^6y^3

OpenStudy (eunice_o):

@Photon336

OpenStudy (eunice_o):

do i take the highest exponent or lowest ?

OpenStudy (ask_if_im_a_ghost):

The lowest

OpenStudy (eunice_o):

when do i ever take the highest exponent ?

OpenStudy (decentnabeel):

\[16v^8 x^9 -28v^2x^6y^3\] \[\mathrm{Apply\:exponent\:rule}:\quad \:a^{b+c}=a^ba\] \[x^9=x^6x^3,\:v^8=v^2v^6\] \[=16x^6x^3v^2v^6-28y^3x^6v^2\] \[\mathrm{Factor\:out\:common\:term\:}4v^2x^6\] \[=4v^2x^6\left(4v^6x^3-7y^3\right)\]

OpenStudy (decentnabeel):

@Eunice_O are you understand

RhondaSommer (rhondasommer):

Hello :D I can help you go through it step by step. would you like that?

OpenStudy (eunice_o):

yes

RhondaSommer (rhondasommer):

okay great. :D give me one second...(i want to solve it so im sure you dont mess up)

RhondaSommer (rhondasommer):

\[What~is~the~largest~number~that~divides~evenly~into~16v^8x^9 and −28v^2x^6y^3?\]

OpenStudy (eunice_o):

4?

RhondaSommer (rhondasommer):

yep :D

OpenStudy (eunice_o):

but how do i know when to take the highest exponent ?

RhondaSommer (rhondasommer):

\[What~is~the~highest~degree~of~v~that~divides~evenly~into\\16v^8x^9~and~−28v^2x^6y^3\]

OpenStudy (eunice_o):

v^8?

RhondaSommer (rhondasommer):

not quite... what is the smallest exponent, 2? or 8?

OpenStudy (eunice_o):

2? but u asked me the highest degree

RhondaSommer (rhondasommer):

that will divide evenly with both of them

RhondaSommer (rhondasommer):

you couldnt divide 8 out of the other one because it only had 2

OpenStudy (eunice_o):

okay but can you please help me with this other question i really do not understand it . Please :)?

RhondaSommer (rhondasommer):

._. do you fully get this one?

OpenStudy (eunice_o):

yes

RhondaSommer (rhondasommer):

post it as a new question